# Shift Parallelism: Low-Latency, High-Throughput LLM Inference for Dynamic Workloads

### v1

@@ Abstract.
We evaluate Shift Parallelism with real-world production traces with dynamic traffic patterns as well as synthetic benchmarking patterns across models, context sizes, and arrival rates. All results affirm the same: Shift Parallelism has a better the latency vs. throughput tradeoff than TP or DP, and hence obtains low latency without degrading throughput in dynamic workloads.
@@ 1. Introduction
LLM inference has become the dominant workload in AI as its applications span agentic systems, chatbox (interactive) applications, model post-training (e.g. reinforcement learning), and image/video generation. The efficiency of inference systems is critical to both the performance and cost of AI applications. As of today, GPU parallelization is the prominent way of enabling large-scale AI production, and advanced multi-GPU parallelization techniques make nontrivial tradeoffs across key performance metrics.
@@
The first two rows of Table 1 shows the performance tradeoffs of TP and DP. We do have the choice to deploy TP and DP in separate nodes and route latency- and throughput-oriented requests, respectively. However, duplicating the node count (one for TP and one for DP) doubles the deployment cost and add complexity.
@@
However, in this work, we notice that Sequence Parallelism (SP) (Jacobs et al., 2023)—another form of parallelism developed and used in training—can offer a potential solution to resolving the challenge. SP splits the input sequence across GPUs to parallelize work within a single request to reduce TTFT. Unlike TP, it avoids costly all-reduce communication, while still achieving high GPU utilization. And while SP cannot parallelize decoding steps, resulting in the worst TPOT compared to TP and DP (Table 1), it has the same KV cache layout as TP, allowing for dynamic switching to TP when TPOT is critical. We call this dynamic approach Shift Parallelism, which is the focus of this work.
@@
Adapt SP from training to inference and generalize it for a diverse set of inference models supporting GQA (Ainslie et al., 2023), load-balancing at small batch sizes, combination of TP, and KV cache replication when paralleism degree is higher than number of KV heads.
Propose Shift Parallelism for dynamically switching across SP and TP for mitigating latency vs througput tradeoff for dynamic workloads.
Test Shift Parallelism with real-world production workloads and evaluate the performance characteristics via extensive benchmarking demonstrating up to 1.5 faster response time with 50% throughput saving.
@@
The rest of the paper is organized as follows. Section 2 provides mode details about production traffic patterns high-performance LLM inference. Section 3 presents SP and Shift Parallelism. Section 4 evaluates the performance of proposed techniques on real-life patterns. Section 5 about related work, and Section 6 concludes the paper.
@@ 2.2. Performance Metrics
Since LLM use cases are diverse, metrics that measure their inference performance is also multi-faceted. In our paper, we focus on three main metrics that cover the most important aspects of interactive and batch workloads:
@@ 2.3. Transformer Architecture
A vanilla LLM involves a series of transformer layers, and each transformer layer consists i) an attention mechanism and a ii) multi-layer perceptron (MLP). The weights in the transformer layer correspond to the QKV (which is a concatenation of q–querry, k–key, and v–value) and O matrices in the attention, as shown in Figure 4.
@@ 2.4. Existing Parallelism Approaches
DP runs multiple replicas across requests and do not accelerate processing of a single request. For accelerating, TP partitions the weight matrices either row-wise or column-wise, as depicted in Figure 3(a). Yet, row parallelization requires all-reduce with communication cost, where is the sequence length. For a fixed sequence length, the communication-to-compute ration increases with the TP degree as shown in the last column of Table 2.
@@
Ulysses has been applied to training, and in this work, we extend Ulysses for inference handing inference specific nuances to allow for a generalized implementation 111In the rest of the paper, we use Ulysses and SP interchangeably..
@@ 3.1. Overview
Figure 5 illustrates KV cache invariance between TP=2 and SP=2. In the center, four attention heads are evenly distributed across two GPUs (two heads per GPU). This distribution is identical under both TP=2 and SP=2, allowing the two configurations to share a single attention mechanism and KV cache.
@@ 3.1.1. SP for Inference
Furthermore, real-world inference is not simply a choice between SP or TP. For optimal performance, systems require arbitrary combinations of SP and TP. Our design supports this flexibility, enabling mixed SP–TP configurations.
@@ 3.1.2. Shift Parallelism
Base configuration: Uses SP or a mixed (SP, TP) setup, as long as SP TP = P, where P is the total number of GPUs.
Shift configuration: Always (SP=1, TP=P), spanning the full node.
@@
Cache invariance: In general, the base and shift configurations are not automatically invariant. For arbitrary SP–TP combinations, head ordering in head parallelism breaks the invariance. We resolve this by developing a general process-to-data mapping to ensure KV cache consistency (3.3.1).
@@ 3.2. SP for Inference
SP is essential for Shift Parallelism because it is the throughput-optimized counterpart of TP as they have the same KV cache layout, and therefore we can switch between two without changing the attention mechanism.
@@ 3.2.1. Design for General Inference
SP is originally implemented for training, and lacks important components, such as GQA mechanism (Ainslie et al., 2023), that is commonly used in inference models. The original MHA mechanism that is described in Section 2.4.
GQA Extension: In this work, we extend SP for GQA mechanism for adapting SP to a diverse set of LLMs that are used in inference. GQA saves memory by sharing each KV head with multiple query heads. However, multi-GPU scaling of GQA is nontrivial with models involving small #KV heads. For example, Qwen-30B-A3B attention has 4 KV heads, and it cannot be scaled to an 8GPU node since there is not enough KV heads to be distributed across more than 4 GPUs.
TP solves this problem by replicating the KV weights in the QKV projection (see Section 2.3), which recomputes the KV cache within GQA groups redundantly. Unfortunately, this solution is not applicable to SP because each process owns only a slice of the input sequence and the missing slices cannot be replicated simply by recomputation.
KV Cache Replication: For SP in inference, we propose an in-network KV cache replication algorithm with multi-step neighborhood collectives. When , i.e., #KV heads, we use a three step algorithm as demonstrated in Figure 6: We i) exchange available KV heads across the GQA groups, resulting neighborhood all-to-all communications within groups . Then ii) we replicate the heads in network using a neighborhood all-gather within GQA groups . As a result of the all-gather, each KV head is replicated within the GQA groups. We finally iii) interleave the incoming sequence to preserve the original sequence order.
Algorithm 1 shows the implementation. Query heads ( uses the SP communicator (Line 3) as usual, and KV heads () use for all-to-all communication (Line 4) and (Line 5) for all all-gather communication. Line 6 shows the reordering that depends on and .
@@
Specifically, decoding in low traffic yields small batch sizes because there is only a few tokens produced at a time. Small batch sizes comparable to the SP degree cannot be evenly partitioned across GPUs, causing serious load imbalance. For example, when the batch size is 9 and , all GPUs to process a single token except the one that processes two tokens, causing 50% efficiency. SP even breaks down when , causing sparse communications.
@@
Fusing Communications: The QKV matrix fuses operations related to q, k, and v, bringing multiple communications down to a single matrix all-to-all communications together as shown in Algorithm 2 Line 4. When KV cache is replicated, the all-to-all symmetry in communications breaks down, and so the fusing. Fusing all the communications in Algorithm 1 communications is not trivial, yet we fuse the all-to-all and all-gather for k and v as a minor optimization.
3.2.2. Combined (SP, TP) Algorithm
We need to combine SP with TP for handling large models that does not fit (or barely fits) in a single GPU. For a throughput-optimal config, we avoid partition the model with TP as much as each partition fits into GPU memory, and there is enough room for KV cache for providing concurrency and high throughput. Then rest of the GPUs can be efficiently employed using SP, which enlarges KV cache. For example, our evaluation involves Llama-17B-16E (FP8) has 109 GB memory footprint, yet needs at least for processing long contexts concurrently in within 141 GB GPU memory, and therefore we need a combination of for an optimal deployment on a node with 8 GPUs.
@@ 3.3. Shift Parallelism
In shift parallelism, we have two configurations; a) the base config that implements SP to optimize TTFT and throughput, and b) the shift configuration that implements full TP to optimize TPOT. The base configuration can optionally be a combination of TP and SP, of model does not fit into a single GPU (Section 3.2.2).
@@
The main criteria of switching between configurations is simple: We choose the base model for large batch size and the shift model for small batch size. Therefore, we decide on a shift parallelism threshold, if the batch size is larger than the threshold, we choose configuration, and choose the shift configuration, i.e., full-TP on group as Algorithm 3 describes.
@@ 3.3.1. General KV Cache Invariance
The KV cache invariance does not only require the same attention head layouts, but also the same ordering of the heads. Interestingly, the invariance across SP and TP breaks down when shifting across arbitrary and configurations. When the base configuration involves a combination of SP and TP, e.g., the attention head order does not follow the same order as in , i.e., not , anymore but . The attention mechanism do not care about the order of the heads as long as we lock in to an order (i.e., base config’s) and stick to it when switching to the shift config as depicted in Figure 7.
Figure 7 shows the distributed memory layout of Q projection with the (a) base config and (b) the shift config. The original config yields TP groups and SP groups . TP groups partition the Q weights across heads (i.e. columns) and replicates the input embeddings. SP groups partitions the embeddings across the sequence (i.e., rows), and replicates the Q weights. As a result of the 2D partitioning, the output of the linear layer () has a global layout as shown in the figure. As a result of the all-to-all communication within the SP groups, the resulting head partitions () results in interleaved head ordering . We need to adjust the attention head ordering of (b) shift config accordingly to provide KV cache consistency.
@@ 3.3.2. Memory Management
There are two ways of implementing Shift Parallelism with generalized KV cache invariance. 1) slicing the model weights on-the-fly and 2) loading separate models that shares the attention mechanism (and the KV cache). We use 2) in our implementation.
@@
Slicing provides the same effect with TP, and has no memory overhead since the running buffer can be reused for all layers. Nevertheless, it is not as performant as the next solution because of each slicing requires matrix transposition due to an FP8 hardware limitation of Hopper tensor cores.
@@
The separate model solution do not share the weights across the base and shift configs, but replicates the weights. In this implementation, we load two separate models, one for the base configuration and one for the shift configuration, and these models share the same KV cache.
When loading the weights for the shift model, we use a separate group, SP_TP that spans both SP and TP groups, but with the order of SP group to preserve KV cache coherency do that the shift model will load the right wieght shards as shown in Figure 7.
@@
where the first and second terms represent the base and shift models’ weights, respectively. As a results, the memory overhead of the shift model is , i.e., an base model with more SP and less TP alleviates the memory overhead of the shift model. For example, when SP=8, the shift model’s memory overhead is 12.5%.
@@ 3.4. Integration into vLLM
For achieving low latency in inference, it is crucial to enable compilation and CUDA graph capture mechanisms in vLLM because the batch size can be small in low traffic. However, capturing all-to-all communications caused compilation issues because their dynamic memory management involved dynamic behaviors that are not supported by vLLM’s compilation infrastructure. As a remedy, we modified compilation by relaxing the assumptions that vLLM make for fully capturing additional GPU kernels and communications related to Shift Parallelism.
@@ 4. Evaluation
In this section, we demonstrate the Shift-Parallelism can mitigate the latency vs throughput tradeoffs commonly seen in TP and DP. More specifically, we show
Shift Parallelism can adapt to bursty real-world traffic pattern, achieving simultaneously lowest latency (up to 3.23 lower) and near optimal throughput compared to TP and DP .
On low traffic Shift Parallelism can achieve up to 6.97 lower TTFT, 2.45 lower TPOT than DP, and up to 51% higher throughput than TP.
@@ 4.1.3. Models
We use the models listed in Table 4, all with FP8 quantization. Shift Parallelism is originally designed for dense models, therefore we first present the main evaluation for L70B and Q32B, and then we discuss the performance limitations with Q30B-3B and L17B-16E which are mixture of experts (MoE) models—their static and active number of parameters are shown separately in Table 4.
@@ 4.1.4. Datasets
We use three three types of datasets. i) synthetic requests with random data for parameterized benchmarking, ii) a trace that is a mixture of requests from HumanEval (Chen et al., 2021) and from a CodeAct agent (Neubig and Wang, 2024) running against SWEBench (Jimenez et al., 2024) (HumenEval is one-shot and SWEBench is agentic), iii) a filtered real-life dataset that matches the synthetic dataset requests for the sake of running speculative decoding, and iv) a bursty traffic pattern that resembles real-life production environment.
@@ 4.2. Latency and Throughput in Real-World Traffic
Table 5 summarizes the latency and throughput statistics that are collected from the trace of the bursty workload experiment (Figure 8). The experiment trace with vLLM’s TP and DP, and also the proposed Shift Parallelism shows that
Shift Parallelism obtains the lowest latency across TP and DP with bursty (dynamic) traffic since can sustain low latency with a higher traffic than the other two .
Shift Parallelism obtains a higher peak throughput than TP, and therefore processes the batches in a shorter time. As a result, there wait time for latency-critical request is reduced significantly, i.e., TTFT does not explode with Shift Parallelism (148 ms vs. 3.9 sec.).
Overall, Shift Parallelism can handle the high-traffic bursts better than both TP ad DP, achieving up to 9.16 lower TTFT and 1.63 lower TPOT that both, while at the same time achiveing nearly as good throughput as DP during high-traffic periods, ultimately, improving the quality of service.

### v2

@@ Abstract.
We evaluate Shift Parallelism with real-world production traces with dynamic traffic patterns as well as synthetic benchmarking patterns across models, context sizes, and arrival rates. All results affirm the same: Shift Parallelism has a better latency vs. throughput tradeoff than TP or DP, and hence obtains low latency without degrading throughput in dynamic workloads.
@@ 1. Introduction
LLM inference has become the dominant workload in AI as its applications span agentic systems, chatbot (interactive) applications, model post-training (e.g. reinforcement learning), and image/video generation. The efficiency of inference systems is critical to both the performance and cost of AI applications. As of today, GPU parallelization is the prominent way of enabling large-scale AI production, and advanced multi-GPU parallelization techniques make nontrivial tradeoffs across key performance metrics.
@@
The first two rows of Table 1 show the performance tradeoffs of TP and DP. We do have the choice to deploy TP and DP in separate nodes and route latency- and throughput-oriented requests, respectively. However, duplicating the node count (one for TP and one for DP) doubles the deployment cost and adds complexity.
@@
However, in this work, we notice that Ulysses Sequence Parallelism (SP) (Jacobs et al., 2023)—another form of parallelism developed and used in training—can offer a potential solution to resolving the challenge. SP splits the input sequence across GPUs to parallelize work within a single request to reduce TTFT. Unlike TP, it avoids costly all-reduce communication, while still achieving high GPU utilization. And while SP cannot parallelize decoding steps, resulting in the worst TPOT compared to TP and DP (Table 1), it has the same KV cache layout as TP, allowing for dynamic switching to TP when TPOT is critical. We call this dynamic approach Shift Parallelism, which is the focus of this work.
@@
Adapt SP from training to inference and generalize it for a diverse set of inference models supporting GQA (Ainslie et al., 2023), load-balancing at small batch sizes, combination of TP, and KV cache replication when parallelism degree is higher than number of KV heads.
Propose Shift Parallelism for dynamically switching across SP and TP for mitigating latency vs throughput tradeoff for dynamic workloads.
Test Shift Parallelism with real-world production workloads and evaluate the performance characteristics via extensive benchmarking demonstrating up to 1.5 faster response time with 50% throughput savings.
@@
The rest of the paper is organized as follows. Section 2 provides more details about production traffic patterns high-performance LLM inference. Section 3 presents SP and Shift Parallelism. Section 4 evaluates the performance of proposed techniques on real-life patterns. Section 5 about related work, and Section 6 concludes the paper.
@@ 2.2. Performance Metrics
Since LLM use cases are diverse, metrics that measure their inference performance are also multi-faceted. In our paper, we focus on three main metrics that cover the most important aspects of interactive and batch workloads:
@@ 2.3. Transformer Architecture
A vanilla LLM involves a series of transformer layers, and each transformer layer consists of: i) an attention mechanism and a ii) multi-layer perceptron (MLP). The weights in the transformer layer correspond to the QKV (which is a concatenation of q—query, k—key, and v—value) and O matrices in the attention, as shown in Figure 4.
@@ 2.4. Existing Parallelism Approaches
DP runs multiple replicas across requests and do not accelerate processing of a single request. For accelerating, TP partitions the weight matrices either row-wise or column-wise, as depicted in Figure 3(a). Yet, row parallelization requires all-reduce with communication cost, where is the sequence length. For a fixed sequence length, the communication-to-compute ratio increases with the TP degree as shown in the last column of Table 2.
@@
Ulysses has been applied to training, and in this work, we extend Ulysses for inference handling inference-specific nuances to allow for a generalized implementation 111In the rest of the paper, we use Ulysses and SP interchangeably..
@@ 3.1. Overview
Figure 5 illustrates KV cache invariance between TP=2 and SP=2. In the center, four attention heads are evenly distributed across two GPUs (two heads per GPU). This distribution is identical under both and , allowing the two configurations to share a single attention mechanism and KV cache.
@@ 3.1.1. SP for Inference
Furthermore, real-world inference is not simply a choice between SP or TP. For optimal performance, systems require arbitrary combinations of SP and TP. Our design supports this flexibility, enabling mixed configurations.
@@ 3.1.2. Shift Parallelism
Base configuration: Uses either full SP or a mixed setup, as long as , where is the total number of GPUs in the node.
Shift configuration: Always , spanning the full node.
@@
Cache invariance: In general, the base and shift configurations are not automatically invariant. For arbitrary combinations, head ordering in head parallelism breaks the invariance. We resolve this by developing a general process-to-data mapping to ensure KV cache consistency in Section 3.3.1.
@@ 3.2. SP for Inference
SP is essential for Shift Parallelism because it is the throughput-optimized counterpart of TP as they have the same KV cache layout, and therefore we can switch between them without changing the attention mechanism.
@@ 3.2.1. Design for General Inference
SP is originally implemented for training, and lacks important components, such as GQA mechanism (Ainslie et al., 2023), that is commonly used in inference models. The original MHA mechanism is described in Section 2.4.
GQA Extension: In this work, we extend SP for GQA mechanism for adapting SP to a diverse set of LLMs that are used in inference. GQA saves memory by sharing each KV head with multiple query heads. However, multi-GPU scaling of GQA is nontrivial with models involving small #KV heads. For example, Qwen-30B-A3B attention has 4 KV heads, and it cannot be scaled to an 8GPU node since there are not enough KV heads to be distributed across more than 4 GPUs.
KV Cache Replication: TP solves the scaling problem by replicating the KV weights in the QKV projection (see Section 2.3), which recomputes the KV cache within GQA groups redundantly. This solution is not applicable to SP because each process owns only a slice of the input sequence and the missing slices cannot be replicated simply by recomputation.
For SP in inference, we implemented the KV cache replication using all-to-all communications. The KV heads are replicated within the send buffers of the collective call, resulting in replication in the receiving buffers across GPUs.
Fusing Communications: The QKV matrix fuses operations related to q, k, and v, bringing multiple communications down to a single matrix all-to-all communications together as represented in Algorithm 1 Line 4.
The GQA implementation replaces with , where is the # q heads and is the # k and v heads, and replicates the KV cache if necessary for handling any combination of and with a single, e.g., fused, all-to-all communications.
@@
Specifically, decoding in low traffic yields small batch sizes because there are only a few tokens produced at a time. Small batch sizes comparable to the SP degree cannot be evenly partitioned across GPUs, causing serious load imbalance. For example, when the batch size is 9 and , all GPUs will process a single token except the one that processes two tokens, causing 50% efficiency. SP even breaks down when , causing sparse communications.
@@
3.2.2. Combined Algorithm
We need to combine SP with TP for handling large models that do not fit (or barely fits) in a single GPU. For a throughput-optimal config, we avoid partitioning the model with TP as much as each partition fits into GPU memory, and there is enough room for KV cache for providing concurrency and high throughput. Then the rest of the GPUs can be efficiently employed using SP, which enlarges KV cache. For example, our evaluation involves Llama-17B-16E (FP8) has 109 GB memory footprint, yet needs at least for processing long contexts concurrently within 141 GB GPU memory, and therefore we need a combination of for an optimal deployment on a node with 8 GPUs.
Algorithm 1 shows the forward pass algorithm with an arbitrary configuration, where is the sequence length, is the hidden dimension, and is the number of heads.
@@ 3.3. Shift Parallelism
In shift parallelism, we have two configurations; a) the base config that implements SP to optimize TTFT and throughput, and b) the shift configuration that implements full TP to optimize TPOT. The base configuration can optionally be a combination of TP and SP (Section 3.2.2), if the model does not fit into a single GPU.
@@
The main criterion of switching between configurations is simple: We choose the base model for large batch size and the shift model for small batch size. Therefore, we decide on a shift parallelism threshold, if the batch size is larger than the threshold, we choose configuration, and choose the shift configuration, i.e., full-TP on group as Algorithm 2 describes.
@@ 3.3.1. General KV Cache Invariance
The KV cache invariance does not only require the same attention head layouts, but also the same ordering of the heads. Interestingly, the invariance across SP and TP breaks down when shifting across arbitrary and configurations. When the base configuration involves a combination of SP and TP, e.g., the attention head order does not follow the same order as in , i.e., not , anymore but . The attention mechanism does not care about the order of the heads as long as we lock in to an order (i.e., base config’s) and stick to it when switching to the shift config as depicted in Figure 6.
Figure 6 shows the distributed memory layout of Q projection with the (a) base config and (b) the shift config. The original config yields TP groups and SP groups . TP groups partition the Q weights across heads (i.e. columns) and replicate the input embeddings. SP groups partition the embeddings across the sequence (i.e., rows), and replicate the Q weights. As a result of the 2D partitioning, the output of the linear layer () has a global layout as shown in the figure. As a result of the all-to-all communication within the SP groups, the head partitions () have interleaved head ordering . We need to adjust the attention head ordering of the shift config accordingly to provide KV cache consistency.
@@ 3.3.2. Memory Management
There are two ways of implementing Shift Parallelism with generalized KV cache invariance. 1) slicing the model weights on-the-fly and 2) loading separate models that share the attention mechanism (and the KV cache). We use 2) in our implementation.
@@
Slicing provides the same effect with TP, and has no memory overhead since the running buffer can be reused for all layers. Nevertheless, it is not as performant as the next solution because each slicing requires matrix transposition due to an FP8 hardware limitation of Hopper tensor cores.
@@
The separate model solution does not share the weights across the base and shift configs, but replicates the weights. In this implementation, we load two separate models, one for the base configuration and one for the shift configuration, and these models share the same KV cache.
When loading the weights for the shift model, we use a separate group, SP_TP that spans both SP and TP groups, but with the order of SP group to preserve KV cache coherency. As a result, the shift model will load the right weight shards as shown in Figure 6.
@@
where the first and second terms represent the base and shift models’ weights, respectively. As a results, the memory overhead of the shift model is , i.e., a base model with more and less alleviates the memory overhead of the shift model. For example, when , the shift model’s memory overhead is 12.5%.
@@ 3.4. Integration into vLLM
For achieving low latency in inference, it is crucial to enable compilation and CUDA graph capture mechanisms in vLLM. The plug-in system compiles and captures both base model and shift model separately. Capturing separate graphs for multiple shapes yields hundreds of graphs, which are registered during initialization and replayed accordingly at runtime. The additional graphs for the shift model do not increase the capturing time or memory significantly.
@@ 4. Evaluation
In this section, we demonstrate that Shift-Parallelism can mitigate the latency vs throughput tradeoffs commonly seen in TP and DP. More specifically, we show
Shift Parallelism can adapt to bursty synthetic traffic pattern, achieving simultaneously lowest latency (up to 3.23 lower) and near optimal throughput compared to TP and DP.
On open-source production traces, Shift Parallelism:
Consistently obtains the lowest TTFT and TPOT, and hence completion time statistics compared to TP and DP when running Azure LLM Code Trace (Patel et al., 2024).
Can keep up with the request traffic with no wait time whereas TP and DP cause growing wait times when running Mooncakce Conversation Trace (Qin et al., 2025).
@@ 4.1.3. Models
We use the models listed in Table 4, all with FP8 quantization. Shift Parallelism is originally designed for dense models, therefore we first present the main evaluation for L70B and Q32B, and then we discuss the performance limitations with mixture of experts (MoE) models—Q30B-A3B and L17B-16E—their static and active number of parameters are shown separately in Table 4.
@@ 4.1.4. Datasets
We use the following datasets in the order of presentation. i) A bursty synthetic pattern that resembles real-life production environment, ii) a couple of open-source production traces (details are given in Section 4.2.2), iii) synthetic requests with random data for parameterized benchmarking, and iv) a trace that is a mixture of requests from HumanEval (Chen et al., 2021) and from a CodeAct agent (Neubig and Wang, 2024) running against SWEBench (Jimenez et al., 2024) (HumenEval is one-shot and SWEBench is agentic). We filter these requests to create a synthetic dataset requests for the sake of running speculative decoding in Section 4.5.
@@ 4.2.1. Bursty Synthetic Workload
Table 5 summarizes the latency and throughput statistics that are collected from the trace of the bursty workload experiment (Figure 7). The experiment trace with vLLM’s TP and DP, and also the proposed Shift Parallelism shows
Shift Parallelism obtains the lowest latency across TP and DP with bursty (dynamic) traffic since it can sustain low latency with a higher traffic than the other two .
Shift Parallelism obtains a higher peak throughput than TP, and therefore processes the batches in a shorter time. As a result, the wait time for latency-critical request is reduced significantly, i.e., TTFT does not explode with Shift Parallelism (148 ms vs. 3.9 sec.).
@@
Overall, Shift Parallelism can handle the high-traffic bursts better than both TP ad DP, achieving up to 9.16 lower TTFT and 1.63 lower TPOT than both, while at the same time achieving nearly as good throughput as DP during high-traffic periods, ultimately, improving the quality of service.
4.2.2. Real-Life Workload Traces
In this section, we study two real-life traces. We performed these case studies on code completion and conversation on a single node based on the following open-source traces: 1) Azure LLM Code Trace (Patel et al., 2024), and 2) Mooncake Conversation Trace (Qin et al., 2025). For demonstration we run both traces for 15 minutes. The resulting request distribution and arrival rates are shown in Figure 8.
Azure LLM Code Trace on Llama-70B
This trace is produced on Azure platform from real-world agentic code generation. We replay these traces on our platform (Section 4.1.1) with the Llama-70B-FP8 model.
As seen in Figure 9, DP (shown with green) is throughput oriented, and handles the bursts better compared to TP (shown with yellow), yielding a lower TTFT and completion time in high traffic. TP is latency oriented and yields a lower TPOT, and hence completion time than DP in low traffic.
Mooncake Conversation Trace on Qwen-32B
This trace is from real-world conversation on Moonshot AI’s platform. In this case, we chose to use a smaller model (Qwen-32B-FP8) because the Llama model did not sustain the traffic and context size in our platform, i.e., the arrival rate is so high for a single node that KV cache becomes full, causing wait times as seen in DP and TP in Figure 10.
In our initial try, Qwen-32B-FP8 did not sustain the conversation traffic. Therefore we turned on FP8 KV cache data type (originally it is FP16) for increasing the KV cache capacity. The wait times were reduced, yet DP and TP still could not sustain the incoming traffic yet Shift Parallelism could run the trace without KV cache overflow. Therefore only SP and Shift Parallel solutions can support this trace on a single node without creating significant queuing delays. The statistics are shown in Figure 11 (b).